If you add 2 mL of 1M HCl to 600 mL of 25 mM phosphate buffer (pH 8.0), what will be the resulting pH? (pKa of phosphate = 7.21)
Answer #1

Total concentration of phosphate = 25 mM

[H2PO4-] + [HPO42-] = 25 mM   ...(1)

According to Henderson-Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log([HPO42-] / [H2PO4-])

8.0 = 7.21 + log([HPO42-] / [H2PO4-])

log([HPO42-] / [H2PO4-]) = 8.0 - 7.21

log([HPO42-] / [H2PO4-]) = 0.79

[HPO42-] / [H2PO4-] = 100.79

[HPO42-] / [H2PO4-] = 6.166 ...(2)

Solving equations (1) and (2), we get

[H2PO4-] = 3.49 mM

[HPO42-] = 21.51 mM

moles H2PO4- = (concentration H2PO4-) * (volume of buffer)

moles H2PO4- = (3.49 mM) * (0.600 L)

moles H2PO4- = 2.1 mmol

moles HPO42- = (concentration HPO42-) * (volume of buffer)

moles HPO42- = (21.51 mM) * (0.600 L)

moles HPO42- = 12.9 mmol

moles HCl added = (concentration HCl) * (volume HCl)

moles HCl added = (1 M) * (2 mL)

moles HCl added = 2 mmol

HCl will convert HPO42- to H2PO4-

new moles H2PO4- = (initial moles H2PO4-) + (moles HCl)

new moles H2PO4- = (2.1 mmol) + (2 mmol)

new moles H2PO4- = 4.1 mmol

new moles HPO42- = (initial moles HPO42-) - (moles HCl)

new moles HPO42- = (12.9 mmol) - (2 mmol)

new moles HPO42- = 10.9 mmol

Again using Henderson-Hasselbalch equation,

pH = pKa + log([HPO42-] / [H2PO4-])

pH = pKa + log(new moles HPO42- / new moles H2PO4-)

pH = 7.21 + log(10.9 mmol / 4.1 mmol)

pH = 7.21 + log(2.65)

pH = 7.21 + 0.42

pH = 7.63

Resulting pH is 7.63